OpenStudy (anonymous):

evaluate function please?? log16X x=1/4

OpenStudy (anonymous):

is 16x, the base of ur log?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

16^y = 1/4 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it seem like negative answer lol

OpenStudy (anonymous):

-1/2 I do believe

OpenStudy (anonymous):

well when u do it out it gives u 16x^ (1/4) = x.

OpenStudy (anonymous):

which then turns into 4 * x^1/4 = x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

canadian math to complex for simple problem :)

OpenStudy (anonymous):

\[\log_{16}x = \frac{1}{4} \]

OpenStudy (anonymous):

my error ..\[\log_{16}\frac{1}{4} = y\]

OpenStudy (anonymous):

y = -1/2

OpenStudy (anonymous):

i thought it was 16x, the base of the logarithm? as in log(16x as base)x=1/4

OpenStudy (anonymous):

what would log10x be when x=10,000 and log10x be when x=4/5??

OpenStudy (anonymous):

common sense say 16x as base is to complicated :)

OpenStudy (anonymous):

uhm actually that would be 12th grade math...where r u from buddy?

OpenStudy (anonymous):

outside of canada

OpenStudy (anonymous):

to answer first one its 4,

OpenStudy (anonymous):

if I give location, house might be gone in morning.

OpenStudy (anonymous):

and the second one is about 6.4.

OpenStudy (anonymous):

lol ok buddy.

OpenStudy (anonymous):

world is cruel when your not canadian ;)

OpenStudy (anonymous):

amy u gotta ask urself in a situation like that, 10 to wat power will give 10 000, and 10 to wat power will give u 4/5

OpenStudy (anonymous):

were not even the bad guys though?...and i know i have family across the world who get treated really badly by their government and pellet,...and if its so cruel come to canada!:P

OpenStudy (anonymous):

10 ^4 will give u 10 000, because 10 x10x10x10 = 10 000

OpenStudy (anonymous):

okay thanks for explaining it!!

OpenStudy (anonymous):

canada have to many hard and fractal raindrops. might have to use gaza healthcare plan if move there

OpenStudy (anonymous):

lol oh boy,